Write the equation of a line that passes through the given point and is perpendicular to the given line. Write the equation in the form of Ax + By=C, where a, b, and c are integers and a>0
(-6,-2);2x-5y=8

Answer by jim_thompson5910(35256)

Anything perpendicular to 2x-5y=8 is of the form 5x+2y = C where C is any number


Note: you swap the x and y coefficients and change the sign of the final y coefficient


Since the line passes through (-6,-2), we know that x = -6 and y = -2


So plug these values into 5x+2y = C to find C


5x+2y = C

5(-6)+2(-2) = C

-30-4 = C

-34 = C

C = -34

So the equation of the perpendicular line is 5x+2y = -34

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim